OpenStudy (anonymous):

1+4/3log little (x-3) 4 = 11/3

OpenStudy (anonymous):

what a pain

OpenStudy (anonymous):

oh not really

OpenStudy (anonymous):

ya i cant figure it out

OpenStudy (anonymous):

1) subtract 1 from both sides get \[\frac{3}{4}\log_{x-3}(4)=\frac{8}{3}\]

OpenStudy (anonymous):

rather \[\frac{4}{3}\log_{x-3}(4)=\frac{8}{3}\]

OpenStudy (anonymous):

multiply both sides by 4/3to get \[ \log_{x-3}(4)=\frac{8}{3}\times \frac{3}{4}=2\]

OpenStudy (anonymous):

rewrite in equivalent exponential form to get \[(x-3)^2=4\] and the solve the quadratic for x

OpenStudy (anonymous):

\[x-3=2\] \[x=5\]

OpenStudy (anonymous):

you can forget about \[x-3=-2\] \[x=1\] because the base of a logarithm cannot be negative

OpenStudy (anonymous):

thanks!